### v1

# Towards Imperceptible Backdoor Attack in Self-supervised Learning

Abstract
Self-supervised learning models are vulnerable to backdoor attacks. Existing backdoor attacks that are effective in self-supervised learning often involve noticeable triggers, like colored patches, which are vulnerable to human inspection. In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models. We first find that existing imperceptible triggers designed for supervised learning are not as effective in compromising self-supervised models. We then identify this ineffectiveness is attributed to the overlap in distributions between the backdoor and augmented samples used in self-supervised learning. Building on this insight, we design an attack using optimized triggers that are disentangled to the augmented transformation in the self-supervised learning, while also remaining imperceptible to human vision. Experiments on five datasets and seven SSL algorithms demonstrate our attack is highly effective and stealthy. It also has strong resistance to existing backdoor defenses. Our code can be found at https://github.com/Zhang-Henry/IMPERATIVE.
1 Introduction
In recent years, Self-Supervised Learning (SSL) has emerged as a powerful paradigm in machine learning, particularly within the computer vision domain. The key advantage of SSL lies in its ability to learn abundant representations from the vast amounts of unlabeled data, bypassing the labor-intensive and costly process of manual labeling. The objective of SSL in computer vision is to develop an image encoder, designed to ensure that similar images yield close embeddings in the encoder. This is achieved by constructing similar images by adding different augmentations on the same image. Utilizing this pre-trained encoder, users can conduct different downstream tasks by training downstream classifiers based on it. The downstream classifiers are typically compact networks with a small number of parameters.
Although SSL has been extensively used in the development of foundational models [5, 22, 4], it is at risk of backdoor attacks [29, 61, 37, 75], where the attacker embeds hidden malicious behavior within the encoder. The backdoor can be inherited to the downstream task. That is, the downstream classifier will predict a specific target label if the input contains a pre-defined backdoor trigger. Existing backdoor attacks on SSL such as BadEncoder [30] achieve high attack success rates. However, a common drawback of these effective attacks is that their trigger patterns are obvious, making them susceptible to human inspection. In this paper, we aim to propose a backdoor attack in SSL that is both effective and stealthy to human vision.
There are various invisible triggers designed for the backdoor attacks on supervised classifiers, such as WaNet [51], ISSBA [41], and filter attack [46]. A straightforward way to achieve imperceptible backdoor attacks in SSL is by directly applying these invisible triggers. However, these existing invisible triggers designed for supervised learning do not perform as well in attacking self-supervised models (Please refer to section 3 and Figure 3 for more details). We then find that this lack of effectiveness is due to the overlapping distributions between the backdoor samples and the augmented samples utilized in SSL. Namely, self-supervised models cannot effectively distinguish the distribution of the backdoor samples and the augmented samples, due to the similarity between the transformation altered by the backdoor trigger and intrinsic image augmentations in SSL, such as RandomGrayscale and ColorJitter (Please refer to section 3 and Figure 3).
Figure 1: Comparison of clean, backdoored samples created by Patch trigger [18], Instagram filter trigger [33], ISSBA trigger [41], WaNet trigger [51] and ours. Their attack success rates tested under the threat model of BadEncoder [30] are beside the name of the methods. Residuals are the difference between clean and backdoored images.
Based on the above observations, we developed a backdoor attack that disentangles its optimized trigger transformation and the augmented transformation in self-supervised learning. In detail, it involves increasing the distributional distance between backdoor samples and the augmented samples in the self-supervised learning process. we also keep the stealthiness of the trigger by adding the constraints on both pixel-space and feature-space distance to the original samples, using metrics like SSIM, PSNR, perceptual loss, and Wasserstein distance. We then implemented our prototype IMPERATIVE (Imperceptible BAckdoor ATtack In Self-superVised LEarning) and tested it on five datasets (i.e., CIFAR10 [34], STL10 [9], GTSRB [68], SVHN [50], ImageNet [60]), and across seven classic SSL frameworks (i.e., SimCLR [5], MoCo [22], BYOL [17], SimSiam [6], DINO [4], SwAV [3] and CLIP [58], please refer to section 5 and Appendix A) with their various augmentation transforms. The results demonstrate that our method is highly effective and stealthy. In detail, it achieves 99.54% average ASR and 0.95 SSIM. As shown in Figure 1, we compare the backdoor residuals and attack success rates of several methods. Our method exhibits the highest attack success rate while maintaining the highest stealthiness. It also effectively bypasses existing backdoor defenses such as DECREE [15], Beatrix [48], STRIP [16], Grad-CAM [64], Neural Cleanse [83], and commonly studied noises such as JPEG compression, Poisson noise, and Salt&Pepper noise.
Our contributions are summarized as follows: ① We observed that existing imperceptible triggers designed for supervised classifiers have limited effectiveness in SSL. ② We find that the reason behind such ineffectiveness is the coupling feature-space distributions for the backdoor samples and augmented samples in the SSL models. ③ Based on our findings, we propose an imperceptible and effective backdoor attack in SSL by disentangling the distribution of backdoor samples and augmented samples in SSL, while constraining the stealthiness of the triggers during the optimization process. ④ Extensive experiments on five datasets and seven SSL algorithms with their different augmentation ways demonstrate our attack is effective and stealthy, and can also be resilient to current SOTA backdoor defense methods.
2 Related Work
2.1 Self-Supervised Learning
The objective of SSL is to utilize a vast array of unlabeled data (referred to as the pre-training dataset) to pre-train an image encoder. After the pre-training phase, this encoder can then be used to develop classifiers for various downstream tasks utilizing a smaller set of labeled data (known as the downstream dataset) [71]. SSL typically includes the following pipelines for contrastive learning as outlined in [27, 19]: ① Negative Examples [5, 22, 98, 77, 78]: This approach fosters closeness among positive examples and seeks to maximize the distance between negative examples within the latent space. Notable methods include SimCLRs [5] and MoCo [22]. ② Self-distillation [4, 6, 17, 78]: This strategy involves the use of two identical deep neural networks, termed siamese networks, which share the same architecture but operate with distinct weights. The aim is to enhance the similarity between two differently augmented versions of the same image under certain conditions, such as BYOL [17], SimSiam [6], and DINO [4]. ③ Clustering [3]: This approach applies a clustering mechanism by employing swapped prediction of the representations derived from both encoders in an end-to-end framework, such as SwAV [3]. We will demonstrate that our method is highly effective and stealthy under the SSL algorithms in each category in section 5 and Appendix A.
2.2 Backdoor Attacks
Numerous strategies for backdoor attacks in SL (supervised-learning) have been introduced [36, 88, 49, 74, 56, 95, 40], which can be classified based on (i) the objectives of the attacks, including attacks tailored to specific inputs [65, 41], specific classes [69], or all inputs [18]; (ii) the methods of attack, such as poisoning the training data [45], identifying susceptible architectures [54], or distributing models with embedded backdoors [28]; and (iii) the criteria for optimization, focusing on the attack’s effectiveness [53], its ability to be transferred [93], or its stealthiness [80, 7, 99, 65]. Since many attacks mentioned above cannot be applied to SSL due to the necessity of labeled data, recent studies have investigated alternative methods for implanting backdoors into models trained with SSL [61, 2, 43, 30]. However, they usually use a patch as a backdoor trigger and are susceptible to human inspection and model simulation.
2.3 Backdoor Defenses
Various defenses have been developed to defend against backdoor attacks [35, 90, 24, 47, 87, 57, 55, 52, 94, 70]. Most of them are designed to protect against backdoor threats targeting supervised classifiers. These mechanisms work by either preventing attacks during the model training process [85, 25, 23, 79], or by identifying and addressing backdoors in already compromised models in an offline manner [73, 100, 72, 86, 92, 66, 46, 39]. Additionally, some methods are aimed at recognizing backdoor-triggered inputs at the time of inference [8, 11, 16]. DECREE [15] and Beatrix [48] are two strategies specifically designed for SSL, and we will demonstrate that our method can bypass them in subsection 5.4.
3 Observations and Insights
Figure 2: Existing imperceptible backdoor triggers, which yield high ASR in supervised learning (SL), do not perform as effectively in self-supervised learning (SSL). The attack framework for SL and SSL are standard backdoor poisoning [18] and BadEncoder [29], respectively.
Figure 3: SimCLR [5] pre-trained model struggled to differentiate between the backdoor samples with WaNet trigger [51] and augmented samples in SimCLR contrastive learning framework.
We first evaluate the effectiveness of existing invisible triggers (i.e., WaNet [51], ISSBA [41], and filter attack [46]) designed for supervised classifiers. In Figure 3, we show their attack success rates on supervised classifiers and self-supervised models. For the experiments on supervised classifiers, the attack method is standard backdoor poisoning [18]. For self-supervised models, we use the attack method introduced in BadEncoder [30] and replace the patch trigger with these invisible triggers. The model used here is ResNet18 [21]. We observed that these invisible triggers, which yield high Attack Success Rates (ASR) in supervised learning, do not perform as effectively in self-supervised learning. In detail, we employed a ResNet18 pre-trained with SimCLR [5] to train and test a binary classification task, aiming to discern whether samples poisoned by WaNet [51] and those augmented through SimCLR’s contrastive learning-based data augmentation are distinguishable. Throughout this process, we maintained consistent feature representations for clean samples between the backdoored and clean models by using utility loss introduced in Jia et al. [32].
Results indicate that the models struggled to differentiate between the two categories. A t-SNE [81] visualization of their features, as presented in Figure 3, indicates a significant overlap and entanglement. From this, we infer that the diminished effectiveness of supervised backdoor attack methods in self-supervised learning scenarios may be attributed to the distributional similarity between the features of contrastive-learning-augmented samples and backdoor samples. Consequently, our method endeavors to disentangle the feature space distribution of these two sets. The experimental results demonstrate that this approach substantially improves ASR in self-supervised learning.
4 Attack Design
4.1 Threat Model
Objectives of the Attacker.  The objective of an attacker is to implant backdoors into a pre-trained image encoder by SSL (self-supervised learning). We define a backdoored image encoder model as $\mathcal{F}_{\theta}$ and the backdoor injector as $\mathcal{I}_{\phi}$. In this way, a downstream classifier trained based on $\mathcal{F}_{\theta}$, which we define as $\mathcal{C}_{\epsilon}$, could produce a specific prediction $c$ designated by the attacker for inputs $\boldsymbol{x}$ implanted with a trigger chosen by the attacker. The formal definition is shown as follows. $y$ here means the correct label of the input $\boldsymbol{x}$.
$\mathcal{C}_{\epsilon}(\mathcal{F}_{\theta}(\boldsymbol{x}))=y,\quad\mathcal{C}_{\epsilon}(\mathcal{F}_{\theta}(\mathcal{I}_{\phi}(\boldsymbol{x})))=c$ (1)
The ultimate aim of the attacker is to modify a clean image encoder and create a backdoored version that accomplishes several key objectives: ① Effectiveness: The backdoored model should achieve a high rate of successful attacks while preserving high accuracy in benign conditions. ② Naturalness and Stealthiness: To avoid detection during human inspection, the sample with the trigger should appear authentic and natural. ③ Robustness: The effectiveness of the attack should remain intact even when the triggered samples undergo common pre-processing steps, such as samples under various noises. Furthermore, the attack should bypass prevailing defense mechanisms.
Attacker’s Knowledge and Capabilities.  We assume that the attacker has access to a clean image encoder pre-trained following [29]. Additionally, it is presumed that the attacker has access to a collection of unlabeled images, referred to as shadow dataset. The attacker is also assumed to have access to a few images from the Internet, called reference inputs, for each combination of a target downstream task and a target class. We assume that the attacker is capable of manipulating the training procedure to create an encoder with embedded backdoors. Accordingly, the attacker also has access to the augmentation transforms used to pre-train the encoder, which can be utilized in IMPERATIVE to generate stealthy and effective backdoor triggers. However, we assume that the attacker cannot interfere with the training process of these downstream classifiers, such as the training dataset, model framework, and weights.
4.2 Preliminaries
HSV and HSL Color Spaces.  HSV and HSL color spaces align more closely with human color perception, unlike RGB’s additive mixing. They allow for intuitive adjustments of tint, shade, and tone through single-dimension modifications. They have wide applications in the computer vision domain [62, 76, 42, 13]. More details are shown in subsection D.5.
Wasserstein Distance.  Wasserstein distance is commonly employed for quantifying differences between two latent distributions without known common support or density functions [71]. Specifically, we focus on the 2-Wasserstein distance, which is elaborated in subsection D.5.
4.3 Overarching Idea
According to the observations in section 3, the key to enhancing the attack success rate in SSL is to disentangle the two overlapping distributions of the backdoor and augmentation transformation in the contrastive learning of the SSL’s pre-training stage. To widen the distributional gap, we initially convert the image from RGB color space to HSV and HSL color spaces, because they align better with the characteristics of augmentative transformations, such as “ColorJitter” (a data augmentation technique that randomly alters the brightness, contrast, saturation, and hue of an image). Since we already know the augmentation ways in the pre-training stage, thus we design $\mathcal{L}_{\text{disentangle}}$ to quantify the distributional gap between images in the two batches, which involves measuring the difference in color characteristics. To further enlarge the distributional gap, we design $\mathcal{L}_{\text{alignment}}$ to pull close the features of backdoor images and reference images.
Moreover, while we try to expand the distributional difference between the backdoor trigger and the augmentation transformation, an excessively large gap might result in a significant divergence between the backdoored image and the original one. This could, in turn, diminish the naturalness of the backdoored image and reduce the stealthiness of the trigger. Hence we design $\mathcal{L}_{\text{stealthy}}$ to blend the backdoor seamlessly with the original image.
4.4 Our Approach: IMPERATIVE
In our context, we refer to a clean pre-trained image encoder and its backdoor-injected one as $\mathcal{F}_{\theta}$ and $\mathcal{F}_{\theta}^{\prime}$. Given any pair of a downstream task and its corresponding target class, labeled as $\left(T_{i},y_{i}\right)$, the attacker gathers a collection of reference inputs denoted by $R_{i}=\{x_{i1},x_{i2},\ldots,x_{ir_{i}}\}$ from the specified target class $y_{i}$, where $r_{i}$ represents the number of reference inputs for $(T_{i},y_{i})$, with $i=1,2,\cdots,t$. Moreover, for each pair $\left(T_{i},y_{i}\right)$, the attacker chooses a trigger $\boldsymbol{e}_{i}$ to implant into samples in the shadow dataset $\mathcal{D}_{s}$. We denote a clean input $\boldsymbol{x}$ embedded with a trigger as $\boldsymbol{x}^{\prime}$, which is called a backdoored input.
4.4.1 Enhancing Distributional Gap Between Backdoor Images and Augmented Images
To effectively enlarge the distributional distinction between the backdoor and the augmentation in SSL, we design $\mathcal{L}_{\text{disentangle}}$ for scenarios where augmentation transformations might weaken or obscure the pattern of backdoor triggers, leading to a decrease in the ASR. Utilizing $\mathcal{L}_{\text{disentangle}}$ ensures that the distinctiveness of the backdoor is maintained even in the face of various image transformations. The disentangle loss is defined as follows.
$\begin{array}[]{rl}\mathcal{L}_{\text{disentangle}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum\limits_{\boldsymbol{x}\in D_{s}}&\|H\left(\boldsymbol{x}^{\prime}\right)-H(\tilde{\boldsymbol{x}})\|_{2}+\|S\left(\boldsymbol{x}^{\prime}\right)-S(\tilde{\boldsymbol{x}})\|_{2}\\ +&\|V\left(\boldsymbol{x}^{\prime}\right)-V(\tilde{\boldsymbol{x}})\|_{2}+\|L\left(\boldsymbol{x}^{\prime}\right)-L(\tilde{\boldsymbol{x}})\|_{2},\end{array}$ (2)
where $H$,$S$,$V$,$L$ denote Hue, Saturation, Value, and Lightness transformations from HSV and HSL color spaces, which are elaborated on in subsection 4.2. We denote an input $\boldsymbol{x}$ augmented by the transformations used in the encoder’s pre-training stage as $\tilde{\boldsymbol{x}}$. $\left|\mathcal{D}_{s}\right|$ denotes the sample number in the shadow dataset. $\|u-v\|_{2}$ denotes the $\ell_{2}$ distance between sample $u$ and sample $v$.
4.4.2 Feature Alignment Between Backdoored and Reference Images
Follwoing BadEncoder [30], we enhance the backdoor attack effectiveness by making the compromised image encoder outputs similar feature embeddings for any sample injected with backdoor $\boldsymbol{x}^{\prime}$ in the shadow dataset $\mathcal{D}_{s}$ and the reference inputs $\mathcal{R}_{i}$ of a pair $\left(T_{i},y_{i}\right)$. Consequently, a compromised downstream classifier developed from our compromised image encoder is inclined to assign identical labels to both reference samples $\mathcal{R}_{i}$ and to any compromised sample $\boldsymbol{x}^{\prime}$. We call this process feature alignment between backdoored and reference images, and the $\mathcal{L}_{alignment}$ is defined as follows.
$\mathcal{L}_{alignment}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}},$ (3)
where $s(\cdot,\cdot)$ is used to quantify the degree of similarity, for instance, cosine similarity, between a pair of feature embeddings. The term $|D_{s}|$ denotes the count of samples within the shadow dataset, and the denominators serve the purpose of standardizing the losses.
4.4.3 Preserving Covert and Natural Backdoors with Expanded Distributional Gaps
We employ several metrics that measure the similarity between the backdoored image and the original one in both pixel and feature space to ensure that our trigger remains both natural and inconspicuous. For assessing similarity in pixel space, we use SSIM [89] and PSNR [26]. Meanwhile, for high-level feature space comparisons, we first use LPIPS [96] which better reflects the subjective experience of image quality and similarity, capturing nuances that simpler metrics might miss. Moreover, since our objective is to reduce the distributional disparity between backdoored and clean samples, the 2-Wasserstein distance (WD) [82] serves as an effective tool for this purpose. We first extract features from backdoored and clean samples respectively using a pre-trained model $\mathcal{M}$, such as ResNet34 [21]. Then we use WD to measure the distance between the extracted features. $\mathcal{L}_{\text{stealthy }}$ is defined as follows.
$\begin{array}[]{rl}\mathcal{L}_{\text{stealthy }}=\sum_{\boldsymbol{x}\in D_{s}}&\lambda_{1}\cdot\mathrm{LPIPS}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)+\mathrm{WD}\left(\mathcal{M}(\boldsymbol{x}^{\prime}),\mathcal{M}(\boldsymbol{x})\right)\\ -&\lambda_{2}\cdot\mathrm{PSNR}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)-\mathrm{SSIM}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right),\end{array}$ (4)
where $\lambda_{1},\lambda_{2}$ are used to scale different loss terms to the same scale from 0 to 1.
Algorithm 1 Our backdoor attack IMPERATIVE
Input: Pre-trained benign encoder $\mathcal{F}_{\theta}^{*}$, shadow dataset $\mathcal{D}_{s}$, reference input set $\mathcal{R}$
      Output: Backdoored encoder $\mathcal{F}_{\theta}$ , backdoor trigger injector $\mathcal{I}_{\phi}$
1:function Ours($\mathcal{F}_{\theta}^{*}$, $\mathcal{D}_{s}$ , $\mathcal{R}$)
2:     $\mathcal{F}_{\theta}\leftarrow\mathcal{F}_{\theta}^{*}$; $\hat{\mathcal{D}_{s}}\leftarrow$ Augment samples in $\mathcal{D}_{s}$;
3:     ϕsubscriptitalic-$\mathcal{I}_{\phi}\leftarrow$ a pre-trained backdoor injector $\mathcal{I}_{\phi}^{*}$ using Algorithm 2
4:     for $\text{iter}=0$ to max_epochs do
5:         ${\mathcal{D}_{s}}^{{}^{\prime}}\leftarrow\mathcal{I}_{\phi}(\mathcal{D}_{s})$
6:         $\mathcal{L}_{\text{disentangle}}\leftarrow$ distribution difference between backdoor images $\boldsymbol{x}^{\prime}$ and augmented images $\tilde{\boldsymbol{x}}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\tilde{\boldsymbol{x}}\in\hat{\mathcal{D}_{s}}$triangleright Equation 2
7:         $\mathcal{L}_{\text{stealthy}}\leftarrow$ distance between backdoor image $\boldsymbol{x}^{\prime}$ and clean image $\boldsymbol{x}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\boldsymbol{x}\in\mathcal{D}_{s}$
8:         $\triangleright$ Equation 4
9:         $\mathcal{L}_{alignment}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}}$triangleright Equation 3
10:         $\mathcal{L}_{\text{injector}}=\mathcal{L}_{\text{stealthy}}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}}$triangleright Equation 5
11:         $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr_{1}\cdot\frac{\partial\mathcal{L}_{\text{injector}}}{\partial\phi_{\mathcal{I}}}$
12:         $\mathcal{L}_{\text{consistency }}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$triangleright Equation 6
13:         $\mathcal{L}_{\text{utility}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$$\triangleright$ Equation 7
14:         $\mathcal{L}_{\text{encoder}}=\mathcal{L}_{\text{alignment}}+\mathcal{L}_{\text{consistency}}+\mathcal{L}_{\text{utility}}$
15:         $\theta_{\mathcal{F}}=\theta_{\mathcal{F}}-lr_{2}\cdot\frac{\partial\mathcal{L}_{\text{encoder}}}{\partial\theta_{\mathcal{F}}}$
4.4.4 Optimization Problem Formulation and Algorithm
We have already defined $\mathcal{L}_{\text{disentangle}}$, $\mathcal{L}_{\text{stealthy }}$, $\mathcal{L}_{alignment}$ in the sections above. Then we can define our IMPERATIVE as an optimization problem. Concretely, our backdoor trigger injector $\mathcal{I}_{\phi}$ is a solution to the subsequent optimization problem:
$\min_{\mathcal{I}_{\phi}}\mathcal{L}=\mathcal{L}_{\text{stealthy }}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}},$ (5)
where $\alpha$ and $\beta$ serve as hyper-parameters to provide equilibrium among these three loss components. We adopt Algorithm 1 to solve the optimization problem, where we alternatively optimize the backdoor injector and the compromised image encoder, and output the final backdoored encoder $\mathcal{F}_{\theta}$ and backdoor trigger injector $\mathcal{I}_{\phi}$. Additionally, to speed up the optimization process and promote the backdoor attack efficacy, we adopt Algorithm 2 to pre-train a backdoor injector to initialize the injector in Algorithm 1 (Please refer to Appendix D.2). We further conduct ablation study to prove the function of the injector pre-training in Appendix B. Finally, we prove that the performance of our model is not sensitive to the hyperparameters in the Appendix C.2.
4.4.5 Highlight of Our Technical Contributions
We wish to highlight our distinct technical contributions: ①Identifying the reason behind existing triggers’ ineffectiveness: We discover that existing backdoor attacks on SSL, such as BadEncoder [30], attain high ASRs but utilize visible triggers. Moreover, traditional imperceptible triggers designed for supervised classifiers, e.g., WaNet [51], ISSBA [41], do not perform effectively in SSL, owing to the overlap in feature-space distributions between the backdoor samples and the augmented samples used in SSL models. ② Proposing disentanglement strategy to generate stealthy and effective backdoor triggers: Based on our insights, we propose a novel backdoor attack in SSL that effectively maintains the potency, stealthiness, and robustness of the generated trigger by disentangling the distribution of backdoor samples from augmented samples within SSL.
5 Evaluation
In this section, we first evaluate the effectiveness and stealthiness of IMPERATIVE using four datasets in subsection 5.2 and subsection 5.3. Then we evaluate the robustness of our method against various SOTA backdoor defenses and noises of different depths in subsection 5.4 and Appendix C.1. After that, we prove our method’s efficacy in large-scale dataset in subsection 5.5. Furthermore, to prove its generalization capability, we conduct more attacking cases on various representative SSL algorithms and a multi-modal model with various augmentations in Appendix A. Moreover, we conduct the ablation study in Appendix B to prove each component’s function. Finally, we evaluate the sensitivity of parameters in Appendix C.2 and further analyze the performance of our method in Appendix C.3.
5.1 Experimental Setup
Datasets.  We utilize the following four image datasets, i.e. CIFAR10 [34], STL10 [9], GTSRB [68], SVHN [50] to evaluate our method, which are also frequently utilized in research related to backdoor attacks [30, 11, 16, 51]. We elaborate on the details of the datasets in Appendix D.6.1.
Pre-training Dataset Downstream Dataset WaNet CTRL Ins-kelvin Ins-xpro2 Ours
[…]
STL10 CIFAR10 86.77 84.43 10.28 87.19 8.72 86.75 18.63 86.85 16.83 87.11 99.58
GTSRB 76.12 74.45 5.23 77.57 8.17 76.49 72.95 76.71 14.02 75.82 97.97
SVHN 55.35 58.29 16.83 54.29 3.32 56.67 38.03 58.42 18.68 58.62 99.76
CIFAR10 STL10 76.14 72.73 9.78 75.73 16.85 74.89 1.16 74.11 5.91 74.48 95.00
GTSRB 81.84 75.85 5.46 79.94 97.95 78.56 2.50 75.08 42.40 79.15 98.73
SVHN 61.52 54.79 17.99 66.33 40.91 68.49 22.13 68.95 30.91 63.67 98.79
Table 1: Results of CA(%), BA(%), and ASR(%) in comparison experiments. Our IMPERATIVE achieves high ASRs while maintaining the accuracy of the downstream classifiers.
Evaluation Metrics.  To assess the effectiveness of our IMPERATIVE, we employ three metrics following existing works [97, 30]: Clean Accuracy (CA), Backdoored Accuracy (BA), Attack Success Rate (ASR). More details are explained in Appendix D.6.2. To evaluate the stealthiness and naturalness of our IMPERATIVE, we employ three metrics following existing works [31, 84]: SSIM [89], PSNR [26] and LPIPS [96]. In experiments, they measure the similarity between backdoored images and original clean images. Higher SSIM and PSNR and lower LIPIPS indicate better stealthiness and naturalness of the generated backdoored images.
Parameter Setting.  We consider a scenario where the attacker chooses a specific downstream task/dataset, a specific target class, and a specific reference input. For the datasets CIFAR10, STL10, GTSRB, and SVHN, the selected target classes are “airplane”, “truck”, “priority sign”, and “digit one”, respectively. CIFAR10 and STL10 are used as the default pre-training dataset. In the pre-training stage, we employ SimCLR [5] by default to train a ResNet18 [21] model, serving as our image encoder. Furthermore, we prove the effectiveness of our method across various representative SSL frameworks in Appendix A, including MoCo [22], BYOL [17], SimSiam [6], DINO [4], SwAV [3], and CLIP [58]. More details are shown in Appendix D.6.3.
Baselines.  We first choose two normal Instagram filters, i.e., kelvin and xpro2, as our baseline triggers, which are gradient-free methods to apply aesthetic enhancements to images, such as adjusting color balance, contrast, and saturation [33, 46]. Moreover, WaNet [51], and CTRL [37], ISSBA [41] are also chosen as our baseline triggers to compare with our method. They are state-of-the-art backdoor triggers and outperform earlier backdoor triggers in terms of stealthiness, yet maintain a high ASR. They are implemented using the default hyperparameters as specified in their original papers. Note that we use BadEncoder [30] to inject these baseline triggers into the compromised encoders.
5.2 Effectiveness Evaluation
Effective Attack.Table 1 indicates our method consistently achieves high ASRs (most above 98%) across various datasets. For instance, with STL10 as the pre-training dataset and CIFAR10 as the downstream dataset, our method reaches an ASR of 99.58%, indicating a highly successful rate of triggering the desired misclassification. This is significantly higher than the ASRs of other benchmark methods like WaNet [51] and CTRL [37] in the same context.
Accuracy Preservation.Table 1 proves that IMPERATIVE preserves the accuracy of the downstream classifiers. The differences between BA and CA are within 2% across the different combinations of pre-training and downstream datasets. This suggests that the presence of the backdoor introduced by our method does not compromise the classifier’s ability to correctly label clean images. For instance, when STL10 is used for pre-training and GTSRB is the downstream dataset, the BA is 76.12% compared to the CA of 75.82%, showing a negligible impact on the classifier’s performance.
[…]
5.3 Stealthiness Evaluation
As shown in Table 3, our method exhibits strong stealthiness advantages with SSIM values consistently close to 1 and most above 0.95, indicating minimal structural changes to the images. PSNR values are notably high, exceeding 40 for some cases, which implies that the noise introduced is hardly detectable. Furthermore, the LPIPS values are exceptionally low, ranging from 0.0001 to 0.02, demonstrating that the perceptual difference between the original and perturbed images is almost negligible. These figures collectively suggest that our method introduces changes that are virtually imperceptible to both human eyes and standard image quality metrics.
5.4 Robustness Evaluation
To assess the resilience of our method against current backdoor defenses, we deploy various state-of-the-art backdoor defense strategies, i.e., DECREE [15], Beatrix [48], Neural Cleanse [83], STRIP [16], Grad-CAM [64] and measure how effectively our method withstands these defenses. Additionally, to further test the robustness of our method, we evaluate its endurance against the following commonly studied noises, i.e., JPEG compression [10, 12, 20, 14], Poisson noise [1, 91], and Salt&Pepper noise [38, 1, 91]. We show the defense results of STRIP, Grad-CAM, and noises in Appendix C.1.
DECREE. [15]  We initially assess whether our method can elude detection by DECREE [15], a method for identifying Trojan attacks in pre-trained encoders. It judges an encoder to be compromised if the proportion of the reversed trigger’s $\mathcal{L}^{1}$ norm relative to the greatest $\mathcal{L}^{1}$ norm within the input space of the model falls below a threshold, i.e., 0.1 $\mathcal{P}\mathcal{L}^{1}$ Norm [15]. As shown in Table 3, $\mathcal{P}\mathcal{L}^{1}$-Norm of each pre-trained and downstream dataset pair are above a default threshold 0.1, and thus this approach does not succeed in identifying backdoored encoders created by IMPERATIVE. This is because the stealthy loss is designed to narrow the distribution gap between backdoored and normal data in both pixel and feature levels. This makes it challenging for DECREE to detect abnormal model internal representation changes when specific behaviors are activated.
Beatrix. [48]  Beatrix [48] identifies poisoned samples by detecting abnormalities in the feature space. We used two pretraining datasets, CIFAR-10 and STL-10, and created backdoored encoders using BadEncoder and IMPERATIVE. By sampling 500 clean inputs and 500 poisoned samples, we applied Beatrix to differentiate them. The findings, as illustrated in Table 5, show that Beatrix effectively recognizes poisoned samples from BadEncoder with over 93% accuracy. However, Beatrix struggles to identify poisoned samples from IMPERATIVE, with a detection accuracy of below 50% on both CIFAR-10 and STL-10, which is like random guessing. We further analyze the reason why Beatrix fails to detect IMPERATIVE in Appendix C.1.1.
[…]
Neural Cleanse. [83]  We then test the resilience of our method against a well-known reverse engineering defense, namely Neural Cleanse (NC) [83]. Specifically, it utilizes the anomaly index to measure the deviation of the reconstructed triggers by their sizes, labeling models with an anomaly index exceeding two as Trojan-infected. Since NC is specifically designed for classifiers and cannot be directly used on image encoders, NC is employed to identify backdoors in a compromised downstream classifier. Moreover, the testing dataset from a targeted downstream dataset serves as the clean dataset. Experimental outcomes, as illustrated in Table 3, indicate that NC is unable to detect the Trojan model created by our approach. While effective in identifying patch-based Trojans [44], this method presumes a uniform trigger pattern at the pixel level across different samples. Our method circumvents this by employing input-dependent triggers, meaning the pixel-level Trojan alterations vary across samples.
5.5 Generalization to Large-scale Dataset
We the generalization ability of our method to large-scale dataset by applying our attack to an image encoder pre-trained on ImageNet [60] released by Google [5]. We show the detailed experimental setup in Appendix D.6.4. Table 5 indicates that our method is highly effective on ImageNet. It achieves high ASR across different datasets, ranging from 96% to 100%, showing its capability to consistently deceive the target system. As shown in Table 3, the high SSIM values ranging from 0.989 to 0.998, high PSNR values ranging from 32.14 to 38.34, and low LPIPS values demonstrate that the perturbations made by IMPERATIVE are almost imperceptible, maintaining the visual quality of the images. Moreover, the BA is close to CA, indicating our attack maintains accuracy for the given downstream task despite the backdoor. Additionally, both our ASR and BA are much higher than those of the benchmark baseline ISSBA [41], proving ours has better performance.
6 Conclusions and Future Work
In this paper, we propose an imperceptible and effective backdoor attack against self-supervised models based on the optimized triggers that is disentangled to the augmented transformation in the self-supervised learning. Based on the evaluation across five different datasets and seven SSL algorithms, our attack is demonstrated to be both highly effective and stealthy. It also effectively bypasses existing backdoor defenses. For future work, it would be beneficial to expand the scope of research to include various other domains of machine learning, such as NLP and audio processing.
Appendix A More Attacking Cases of IMPERATIVE
A.1 Generalization to Other Representative SSL Frameworks
As elaborated on subsection 2.1, SSL frameworks for Contrastive Learning can be normally based on the following categories: negative examples, self-distillation, and clustering. To further evaluate the efficacy of IMPERATIVE, we test our method across the representative SSL algorithms in each category apart from the default SimCLR, i.e., negative examples (MoCo [22]), self-distillation (BYOL [17], SimSiam [6], and DINO [4]), and clustering (SwAV [3]). Table 7 illustrates the detailed parameters for different augmentation transforms used in the SSL algorithms, including Random resized crop, Random horizontal flip, Color jitter, Random gray scale, and Random polarization. The backbone feature extractor used here is ResNet18 [21]. Other experiment settings are the same as those in Appendix D.6.3.
Table 6 shows the evaluation results of applying IMPERATIVE to the encoders pre-trained utilizing various SSL algorithms on CIFAR-10. The SSL algorithms demonstrate high effectiveness in achieving successful backdoor attacks with ASR nearing or reaching perfection in many cases, with an average ASR of 99.54%. Additionally, the algorithms excel in maintaining the stealthiness of these attacks, evidenced by a high average SSIM of 0.95 and a high average PSNR value of 26.89, alongside a low LPIPS value of 0.011. These metrics across the SSL algorithms collectively indicate that while IMPERATIVE are highly effective at manipulating model behaviors, they do so in a way that remains largely undetectable through standard image quality assessments, which aligns with the conclusions drawn in section 5.
[…]
MOCO [22] STL10 72.56 99.75 0.95 26.74 0.0080
GTSRB 68.11 99.98 0.95 26.84 0.0070
SVHN 74.58 99.97 0.96 31.86 0.0030
SimSiam [6] STL10 57.56 100.00 0.95 23.29 0.0230
GTSRB 28.99 99.90 0.96 24.26 0.0190
SVHN 52.24 100.00 0.96 26.04 0.0141
BYOL [17] STL10 39.31 98.33 0.95 26.94 0.0070
GTSRB 11.15 99.31 0.96 32.35 0.0020
SVHN 42.76 99.74 0.96 31.88 0.0024
SwAV [3] STL10 66.28 99.76 0.89 23.43 0.0258
GTSRB 79.39 100.00 0.97 25.42 0.0089
SVHN 76.71 96.94 0.96 25.05 0.0096
DINO [4] STL10 36.06 99.57 0.93 26.98 0.0141
GTSRB 11.49 99.83 0.95 27.11 0.0069
SVHN 37.66 99.99 0.93 25.18 0.0184
Average 50.32 99.54 0.95 26.89 0.0110
Table 6: Evaluations on other SSL algorithms, which are used to pre-train the encoders on CIFAR-10. IMPERATIVE keeps highly effective and stealthy across these algorithms.
Parameter SimCLR/CLIP MOCO SimSiam BYOL SwAV DINO Description
input_size 32/224 32 32 32 / / Size of the input image in pixels.
cj_prob 0.8 0.8 0.8 0.8 0.8 0.8 Probability that color jitter is applied.
cj_strength 1 0.5 1 1 0.5 0.5 Strength of the color jitter. cj_bright, cj_contrast, cj_sat, and cj_hue are multiplied by this value.
cj_bright 0.4 0.4 0.4 0.4 0.8 0.8 How much to jitter brightness.
cj_contrast 0.4 0.4 0.4 0.4 0.8 0.8 How much to jitter contrast.
cj_sat 0.4 0.4 0.4 0.2 0.8 0.4 How much to jitter saturation.
cj_hue 0.1 0.1 0.1 0.1 0.2 0.2 How much to jitter hue.
min_scale / 0.2 0.2 0.08 / / Minimum size of the randomized crop relative to the input_size.
random_gray_scale 0.2 0.2 0.2 0.2 0.2 0.2 Probability of conversion to grayscale.
vf_prob 0 0 0 0 0 0 Probability that vertical flip is applied.
hf_prob 0.5 0.5 0.5 0.5 0.5 0.5 Probability that horizontal flip is applied.
rr_prob 0 0 0 0 0 0 Probability that random rotation is applied.
crop_sizes / / / / [32,32] / Size of the input image in pixels for each crop category.
crop_counts / / / / [2,2] / Number of crops for each crop category.
crop_min_scales / / / / [0.14,0.14] / Min scales for each crop category.
crop_max_scales / / / / [32,32] / Max_scales for each crop category.
global_crop_size / / / / / 32 Crop size of the global views.
global_crop_scale / / / / / [0.4, 1.0] Tuple of min and max scales relative to global_crop_size.
local_crop_size / / / / / 96 Crop size of the local views.
local_crop_scale / / / / / [0.05, 0.4] Tuple of min and max scales relative to local_crop_size.
n_local_views / / / / / 6 Number of generated local views.
solarization_prob / / / / / 0.2 Probability to apply solarization on the second global view.
Table 7: Parameters of Augmentation Transforms in SSL Algorithms.
A.2 Generalization to Multi-modal Models
We utilize our IMPERATIVE to integrate a backdoor into the image encoder of the multi-modal model CLIP [58], which is comprised of both an image and a text encoder, pre-trained on 400 million internet-sourced (image, text) pairs. This backdoored image encoder is then employed to construct a multi-shot downstream classifier.
Due to the unavailability of CLIP’s original pre-training dataset, we utilize the training images from CIFAR-10, resized to 224×224×3, as our shadow dataset. We define $\alpha=0.1$ and $\beta=20$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to normalize the loss terms across similar scales. Additionally, Algorithm 1 and Algorithm 2 are configured with a batch size of 16, and learning rates are set at $10^{-6}$ and 0.005, respectively, for a duration of 100 epochs. SVHN serves as the downstream dataset with the target label “one”. We employ the same reference inputs as described in [29]. All other experimental settings remain consistent with those outlined in Appendix D.6.3.
Table 8 presents the attack results of multi-shot classification using CLIP. IMPERATIVE achieves a similar high BA and ASR as BadEncoder [29] does, but it maintains much more stealthiness with high SSIM&PSNR and low LPIPIS values than BadEncoder which uses a white patch as the backdoor trigger.
Methods CA(%) BA(%) ASR(%) SSIM PSNR LPIPIS
Ours 70.6 68.98 99.95 0.98 30.11 0.0179
BadEncoder [29] 70.27 99.99 / / /
Table 8: Attack results on the multi-shot classification of CLIP. IMPERATIVE attains comparable high BA and ASR to those of BadEncoder [29], yet it offers significantly enhanced stealthiness.
Appendix B Ablation Study
As discussed in subsection 4.4, we utilize $\mathcal{L}_{disentangle}$ to enlarge the gap between the backdoored image distribution and the augmented image distribution. We use $\mathcal{L}_{alignment}$ to promote feature alignment between backdoored images and reference images. We use $\mathcal{L}_{stealthy}$ to make the backdoor trigger stealthy and natural. In Algorithm 1, we initialize a backdoor injector using a pre-trained one by Algorithm 2. We use STL10 as the dataset for pre-training encoders. In this section, we demonstrate the effects of each important component in subsection 4.4 by ablating them respectively.
As reported in Table 9, the ASRs of our method decrease across all target downstream datasets when ablating each component respectively. As reported in Table 10, the average SSIM and PSNR drop dramatically and LPIPS increases a lot across all datasets when ablating $\mathcal{L}_{stealthy}$. For example, in the ablation case of the target downstream dataset CIFAR10, the ASR decreases from 99.58% to 67.84%, 10.22%, 75.27% when ablating $\mathcal{L}_{disentangle}$, $\mathcal{L}_{alignment}$ and backdoor injector initialization respectively. This distinctly shows that the absence of key components significantly compromises the effectiveness of our backdoor attack methodology. Moreover, when ablating $\mathcal{L}_{stealthy}$, the SSIM drops from 0.99 to 0.1. The PSNR drops from 41.80 to 5.65. The LPIPS increases from 0.0002 to 0.55. This indicates that the backdoored images without the $\mathcal{L}_{stealthy}$ are more easily detectable and differ a lot from the original images.
Pre-training Backdoor injector $\mathcal{L}_{\text{disentangle}}$ $\mathcal{L}_{\text{alignment}}$ $\mathcal{L}_{\text{stealthy }}$ Target Downstream Dataset
CIFAR10 GTSRB SVHN
[…]
84.66 75.27 78.72 73.33 58.38 13.87
85.71 67.84 74.05 52.26 56.41 62.56
87.20 10.22 65.49 7.46 56.01 23.44
87.11 99.58 75.82 97.97 58.62 99.76
Table 9: ASR(%), BA(%) in ablation studies. Ablating the $\mathcal{L}_{\text{alignment}}$ means ablating the one in $\mathcal{L}_{\text{injector}}$ instead of $\mathcal{L}_{encoder}$, and ablating the pre-trained backdoor injector means randomly initializing it.
Method Target Downstream Dataset
CIFAR10 GTSRB SVHN
[…]
w/o $\mathcal{L}_{stealthy}$ 0.10 5.65 0.5500 0.02 5.20 0.4400 0.10 4.85 0.52
Ours 0.99 41.80 0.0002 0.99 43.58 0.0001 0.96 25.65 0.01
Table 10: Results of ablating $\mathcal{L}_{stealthy}$. Images backdoored without the $\mathcal{L}_{stealthy}$ are more readily identifiable and exhibit significant deviations from the original images.
Appendix C More Evaluations
C.1 More Robustness Tests
C.1.1 Beatrix
Further examination of Beatrix’s detection performance reveals that it determines whether a sample is poisoned based on the deviation in the feature matrix within the embedding space [48]. A sample is deemed poisoned if the deviation exceeds a specific threshold. Figure 4 displays the distribution of deviation values. Notably, poisoned samples by BadEncoder show significantly higher deviation than clean samples, enabling Beatrix to achieve high detection accuracy. Conversely, the deviation distributions between clean inputs and poisoned samples by IMPERATIVE overlap considerably, making them difficult to differentiate.
Figure 4: Beatrix’s deviation distribution [48] for clean and backdoored data, where blue bars represent clean samples and orange bars indicate poisoned ones. Poisoned samples from BadEncoder display markedly higher deviations than clean samples, allowing Beatrix to detect them effectively. However, the deviation distributions for clean inputs and poisoned samples from IMPERATIVE significantly overlap, complicating their differentiation.
Figure 5: Resilience to Grad-CAM [64]. The resemblance observed in these heatmaps indicates that IMPERATIVE is capable of resisting defenses that rely on Grad-CAM.
C.1.2 Grad-CAM
We assess IMPERATIVE’s resilience to Grad-CAM [64], which generates a heatmap for a given model and input sample, where the heat intensity of each pixel reflects its significance in the model’s final prediction. Grad-CAM is effective for identifying smaller Trojans [45, 18], as these Trojans tend to produce high heat values in compact trigger zones, leading to abnormal heatmaps. Nevertheless, our backdoor transformation function alters the entire image, rendering Grad-CAM ineffective in detecting it. Figure 5 illustrates the comparison of heatmaps from both clean images and backdoored images created by our method. The similarity in these heatmaps suggests that IMPERATIVE can withstand defenses based on Grad-CAM.
C.1.3 Various Noises
Next, we are going to introduce the noises, i.e., JPEG compression, Salt&Pepper noise, and Poisson noise, applied in the robustness test.
  • JPEG Compression: JPEG is a prevalent image format. It is a common case that images undergo compression, especially during web transmission. The quality scale ranges from 1 to 100, with 100 being the highest quality and 1 being the lowest quality.
  • Poisson Noise: It arises from the statistical nature of photon reception by image sensors, resulting in random noise of varying intensities. It is set between 1 and 10 to model different lighting conditions and sensor sensitivities.
  • Salt&Pepper Noise: This noise model mimics random pixel corruption due to image sensor errors, transmission faults, or system failures, characterized by randomly scattered white (salt) and black (pepper) pixels. It typically ranges from 0 to 1, but practical applications often use a much smaller range (e.g., 0.01 to 0.2) to avoid overwhelming the image with noise.
Figure 6: Resilience to JPEG compression, Salt&Pepper noise, Poisson noise of different densities.
Evaluating the robustness of our method across these noises helps ensure its performance is reliable even when image quality is compromised when encountering real-world image issues. We illustrate the evaluation results in Figure 6, where the pre-trained dataset is CIFAR10 and the downstream dataset is STL10. The results prove that our method, which utilizes $\mathcal{L}_{disentangle}$, maintains high ASRs in most cases, and the ASR is significantly higher than the cases ablating $\mathcal{L}_{disentangle}$ in our method. For example, under different intensities of Poisson noise, the ASR of our method maintains nearly 100%, while the ASR of the cases ablating $\mathcal{L}_{disentangle}$ in our method is unstable, ranging from 12% to 61%. The outcome indicates that the $\mathcal{L}_{disentangle}$ plays a prominent role in promoting the robustness of our method. The reason may be that the $\mathcal{L}_{disentangle}$ encourages the model to learn features that remain invariant under the transformations. Namely, the model can still recognize the embedded triggers even when various noises or other image perturbations are applied. This aspect of feature preservation is key to ensuring that the backdoor attack remains effective regardless of the image transformations or certain noises applied.
C.1.4 STRIP
Figure 7: Resilient to STRIP [16]. IMPERATIVE can withstand the STRIP runtime defense.
STRIP [16] scrutinizes a suspect sample by overlaying diverse image patterns onto it and monitoring the uniformity of the model’s predictions for these altered inputs. A low entropy score, suggesting consistent predictions across these perturbed samples, would lead STRIP to classify the sample as Trojan-infected. The experimental outcomes depicted in Figure 7 reveal that the entropy values for both clean and Trojan-compromised models produced through our method overlap considerably, signifying that our attack can withstand the STRIP runtime defense. IMPERATIVE manages to evade STRIP as the superimposition process destroys the attack’s trigger, thereby causing both the Trojan-induced and clean samples to exhibit significant alterations in prediction when superimposed, aligning them with what is observed in clean cases.
C.2 Sensitivity Evaluation
Figure 8: Sensitivity evaluation of hyperparameters. IMPERATIVE performs consistently over these parameter ranges.
Figure 8 illustrates the sensitivity evaluation of the hyperparameters alpha and beta, showcasing that the model’s performance exhibits relative insensitivity to parameter variations. Specifically, for alpha, both the Attack Success Rate (ASR) and Backdoored Accuracy (BA) remain relatively constant across a range of alpha values from 0.05 to 0.30, suggesting that once an optimal threshold is reached, the model’s performance is stable despite further adjustments to alpha. Similarly, for beta, changes across a broader spectrum, from 0 to 10, show that ASR quickly stabilizes at 100%, while BA and CA display minimal fluctuations, indicating the robust model performance that is not significantly affected by variations in beta. The horizontal axis in both graphs represents the incremental values of alpha and beta, demonstrating the model’s consistent performance over these parameter ranges.
C.3 Results Analysis
Figure 9: Visualization of feature space segregation over epochs demonstrates the efficacy of our backdoor trigger in creating a distinguishable separation between backdoor and clean features, unlike WaNet [51] where the features remain intermingled, highlighting our method’s utility in securing encoders against backdoor threats.
Figure 9 presents a visual analysis of the feature shifts induced by backdoor injection across different epochs for our proposed method compared to the WaNet [51]. Throughout the backdoor injection process, it becomes evident that our method’s trigger effectively separates the backdoor features from the clean features. This distinction is observable as the training progresses from epoch 10 through to epoch 200, where the separation between the backdoor and clean features becomes increasingly pronounced, indicating a clear demarcation that our encoder can exploit to differentiate between the two.
In contrast, the bottom row representing WaNet’s performance shows a significant overlap of backdoor and clean features. Even at epoch 200, the features remain intermixed, demonstrating WaNet’s inability to distinguish effectively between the two, as the clean, backdoor, and reference features are all entangled. The effectiveness of our backdoor trigger is thus underscored. It introduces a discernible feature shift that an encoder can leverage to recognize backdoored inputs, validating the practical utility of our approach in enhancing model security against backdoor attacks.
Appendix D Implementation Details
The following sections provide a comprehensive breakdown of the process of implementing the attack, which is methodically organized into three distinct phases: pre-training the encoder (Appendix D.1), pre-training the backdoor injector (Appendix D.2), injecting the backdoor (Appendix D.3), and training the downstream classifiers (Appendix D.4). Additionally, we explain the HSV&HSL Color Spaces and Wasserstein Distance (Appendix D.5) used in the method in detail. Moreover, we provide more details on the datasets (Appendix D.6.1), evaluation metrics (Appendix D.6.2), and experimental setup details (Appendix D.6.3, Appendix D.6.4).
D.1 Pre-training Image Encoders
The initial step involves pre-training an image encoder using a specific dataset, hereinafter referred to as the pre-training dataset. In alignment with the experimental framework established in BadEncoder [30], we opt for CIFAR10 or STL10 as our pre-training datasets. These choices are informed by their relatively large size and superior data quality. Crucially, our experimentation leverages the pre-trained weights from BadEncoder’s image encoders, which underwent training over 1,000 epochs employing the Adam optimizer at a learning rate of 0.001 [29]. Specifically, for the CIFAR10 dataset, the encoder is pre-trained using only the training images, excluding their labels. Meanwhile, for STL10, the pre-training also incorporates its additional unlabeled images.
D.2 Pre-training Backdoor Injector
To enhance the optimization speed and increase the effectiveness of the backdoor attack, we use the method outlined in Algorithm 2 to pre-train a backdoor injector, which serves as the initial setup for the injector in Algorithm 1. We use the U-Net architecture [59] for the backdoor injector. The detailed architecture is shown in the Table 11.
Algorithm 2 Pre-training backdoor injector
Input: Shadow dataset $\mathcal{D}_{s}$       Output: Pre-trained backdoor injector $\mathcal{I}_{\phi}$
1:function Pre-training injector($\mathcal{D}_{s}$)
2:     $\hat{\mathcal{D}_{s}}\leftarrow$ Augment samples in $\mathcal{D}_{s}$; $\mathcal{I}_{\phi}\leftarrow\text{Random initialization}$
3:     for $\text{iter}=0$ to max_epochs do
4:         ${\mathcal{D}_{s}}^{{}^{\prime}}\leftarrow\mathcal{I}_{\phi}(\mathcal{D}_{s})$
5:         $\mathcal{L}_{\text{disentangle}}\leftarrow$ distribution difference between backdoor images $\boldsymbol{x}^{\prime}$ and augmented images $\tilde{\boldsymbol{x}}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\tilde{\boldsymbol{x}}\in\hat{\mathcal{D}_{s}}$triangleright Equation 2
6:         $\mathcal{L}_{\text{stealthy}}\leftarrow$ distance between backdoor image $\boldsymbol{x}^{\prime}$ and clean image $\boldsymbol{x}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\boldsymbol{x}\in\mathcal{D}_{s}$
7:         $\triangleright$ Equation 4
8:         $\mathcal{L}_{\text{ours}}=\mathcal{L}_{\text{stealthy}}+\mu\cdot\mathcal{L}_{\text{disentangle}}$
9:         $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr\cdot\frac{\partial\mathcal{L}_{\text{ours}}}{\partial\phi_{\mathcal{I}}}$
Module Layer Type Kernel Size Stride Padding Channel I/O Normalization Activation Input
Encoder Maxpool MaxPool2d 2x2 2 0 - - - Input Image
Conv1 Conv2d 3x3 1 1 3/8 BatchNorm2d (8) ReLU Maxpool
Conv2 Conv2d 3x3 1 1 8/16 BatchNorm2d (16) ReLU Conv1
Conv3 Conv2d 3x3 1 1 16/32 BatchNorm2d (32) ReLU Conv2
Conv4 Conv2d 3x3 1 1 32/64 BatchNorm2d (64) ReLU Conv3
Conv5 Conv2d 3x3 1 1 64/128 BatchNorm2d (128) ReLU Conv4
Decoder Up5 Upsample + Conv2d -/3x3 -/1 -/1 128/64 BatchNorm2d (64) ReLU Conv5
Up_conv5 Conv2d 3x3 1 1 128/64 BatchNorm2d (64) ReLU Up5
Up4 Upsample + Conv2d -/3x3 -/1 -/1 64/32 BatchNorm2d (32) ReLU Up_conv5
Up_conv4 Conv2d 3x3 1 1 64/32 BatchNorm2d (32) ReLU Up4
Up3 Upsample + Conv2d -/3x3 -/1 -/1 32/16 BatchNorm2d (16) ReLU Up_conv4
Up_conv3 Conv2d 3x3 1 1 32/16 BatchNorm2d (16) ReLU Up3
Up2 Upsample + Conv2d -/3x3 -/1 -/1 16/8 BatchNorm2d (8) ReLU Up_conv3
Up_conv2 Conv2d 3x3 1 1 16/8 BatchNorm2d (8) ReLU Up2
Conv_1x1 Conv2d 1x1 1 - 8/3 - - Up_conv2
Table 11: The detailed network structure of backdoor injector.
D.3 Backdoor Injection
The subsequent phase introduces a backdoor into the pre-trained encoder. This is achieved through the use of Shadow Dataset, which constitutes a smaller subset of the pre-training dataset. Consistent with preceding studies [29], this shadow dataset comprises 50,000 randomly selected training samples. Additionally, we employ the same reference samples as those used in BadEncoder [29]. The backdoor is injected using a designated algorithm, which leverages a combination of $\mathcal{L}_{consistency}$, $\mathcal{L}_{utility}$, and $\mathcal{L}_{alignment}$. The notations are the same as the main paper.
Consistency Loss. The consistency loss decreases when the feature vectors generated by the backdoored image encoder and the clean image encoder for the reference inputs are more alike.
$\mathcal{L}_{consistency}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$ (6)
Utility Loss. Utility loss diminishes when the backdoored image encoder and the clean image encoder yield more closely matching feature vectors for a clean input within the shadow dataset.
$\mathcal{L}_{utility}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$ (7)
D.4 Training Downstream Classifiers
Finally, using the pre-trained image encoder, we proceed to train classifiers for three separate datasets, hereafter referred to as downstream datasets. Our approach utilizes a two-layer fully connected neural network as the classifier architecture, with the first and second layers consisting of 512 and 256 neurons, respectively. The testing subset of each downstream dataset is utilized for evaluation purposes. The training of the classifier is conducted over 500 epochs using the cross-entropy loss function and the Adam optimizer, with a set learning rate of 0.0001.
D.5 Technical Background
HSV Color Space. The HSV color model describes color based on three primary attributes: hue, saturation, and value. Hue (H) represents the color itself, essentially the aspect of color we typically refer to by names like red or yellow. Saturation (S) indicates the vividness or richness of the color. A higher saturation means a more intense, pure color. Value (V), often referred to as brightness, measures the lightness or darkness of the color [67]. HSV color space can be converted from RGB color space through Equation 8, Equation 9, Equation 10, Equation 11 [63].
HSL Color Space. The HSL color model is defined by three cylindrical coordinates: hue (H), saturation (S), and lightness (L), which together capture the subtleties of color. HSL color space can be converted from RGB color space through Equation 8, Equation 9, Equation 12, Equation 13 [63].
$\mathrm{R}^{\prime}=\mathrm{R}/255;$ (8)
$\mathrm{G}^{\prime}=\mathrm{G}/255;$
$\mathrm{B}^{\prime}=\mathrm{B}/255$
$\mathrm{C}_{\max}=\operatorname{MAX}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right)$
$\mathrm{C}_{\min}=\operatorname{MIN}\left(\mathrm{R}^{\prime},\mathrm{G}^{\prime},\mathrm{B}^{\prime}\right)$
$\Delta=\mathrm{C}_{\max}-\mathrm{C}_{\min}$
$\mathrm{H}=\left\{\begin{array}[]{l}60^{\circ}\times\left(\frac{\mathrm{G}^{\prime}-\mathrm{B}^{\prime}}{\Delta}\bmod 6\right),\mathrm{C}_{\max}=\mathrm{R}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{B}^{\prime}-\mathrm{R}^{\prime}}{\Delta}+2\right),\mathrm{C}_{\max}=\mathrm{G}^{\prime}\\ 60^{\circ}\times\left(\frac{\mathrm{R}^{\prime}-\mathrm{G}^{\prime}}{\Delta}+4\right),\mathrm{C}_{\max}=\mathrm{B}^{\prime}\end{array}\right.$ (9)
$\mathrm{S}=\left\{\begin{array}[]{cc}0\quad,&\Delta=0\\ \frac{\Delta}{\mathrm{C}_{\max}},&\Delta\neq 0\end{array}\right.$ (10)
$\mathrm{V}=\mathrm{C}_{\max}$ (11)
$\mathrm{S}=\left\{\begin{array}[]{cl}0,&\Delta=0\\ \frac{\Delta}{1-|2\mathrm{L}-1|}&,\Delta\neq 0\end{array}\right.$ (12)
$\mathrm{L}=\left(\mathrm{C}_{\max}+\mathrm{C}_{\min}\right)/2$ (13)
Wasserstein Distance.  Wasserstein distance is commonly employed for quantifying differences between two latent distributions without known common support or density functions [71]. Specifically, we focus on the 2-Wasserstein distance, which is defined below.
[…]
where $\xi$ and $\tau$ correspond to the marginal probability distributions of the clean and poisoned datasets. $\psi$ denotes the combined distribution of both clean and poisoned data samples. $\inf$ represents the infimum, which is the lowest boundary value for the calculated distances within the scope of the joint distribution $\psi$. The integral part of the equation aggregates the distances between each pair of data points, $(u,v)$, where one belongs to the clean set and the other to the poisoned set, as drawn from the joint distribution $\psi$. Here, $p(u,v)$ defines the probability of concurrently selecting both data samples. The term $||u-v||_{2}$ measures the $L^{2}$ distance between the two data points [82].
D.6 Experiments Details
D.6.1 Dataset Details
We use the following datasets in our method evaluation.
  • CIFAR10 [34]: This dataset comprises 60,000 images of 32×32×3 pixels and 10 different classes for basic image recognition tasks, divided into 50,000 for training and 10,000 for testing.
  • STL10 [9]: STL10 includes 5,000 labeled training images and 8,000 for testing with a resolution of 96×96×3 pixels, across 10 classes. Additionally, it provides 100,000 unlabeled images for unsupervised learning. Notably, they are resized to 32×32×3 to be consistent with other datasets.
  • GTSRB [68]: This dataset includes 51,800 images of traffic signs categorized into 43 classes. It is split into 39,200 training and 12,600 testing images, each sized at 32×32×3.
  • SVHN [50]: SVHN is a dataset of digit images from house numbers in Google Street View, consisting of 73,257 training and 26,032 testing images, each 32×32×3 in size.
  • ImageNet [60]: ImageNet is designed for large-scale object classification, featuring 1,281,167 training samples and 50,000 testing samples across 1000 categories. Each image has a resolution of 224x224 pixels with three color channels.
D.6.2 Evaluation Metrics Details
The meanings of evaluation metric are shown below:
  • CA: This is the accuracy of a clean downstream classifier in correctly classifying the clean testing images from the corresponding downstream dataset.
  • BA: BA measures the accuracy of a backdoored downstream classifier on the same clean testing images from the corresponding downstream dataset.
  • ASR: ASR measures the percentage of these backdoored test images that are classified as the target class by the backdoored downstream classifier.
D.6.3 Parameter Setting Details
We adopt the default parameter settings as follows unless stated otherwise:
  1. 1.
    We set $\lambda_{1}=10,\lambda_{2}=0.025$ in the loss equation term Equation 4, $\alpha=0.1,\beta=5$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to scale the loss terms to the same 0-1 range.
  2. 2.
    The Shadow dataset consists of 50000 images randomly chosen from the pre-training dataset.
  3. 3.
    Cosine similarity is adopted to measure the similarity between two samples’ feature embeddings outputted by an encoder.
  4. 4.
    Since we know the SSL’s pre-training process, the SSL augmentation transforms are also known. By default, the augmentation transforms used are RandomResizeCrop, RandomHorizontalFlip, ColorJitter, and RandomGrayscale.
  5. 5.
    We adopt U-Net [59] as the structure of the backdoor injector, which is widely used in image-to-image tasks.
  6. 6.
    Algorithm 1 and Algorithm 2 utilize the Adam optimizer with 200 training epochs, a batch size of 256, and learning rates of 0.001 and 0.005 respectively.
D.6.4 Experimental Setup in Real World Case ImageNet
We selected a random 1% subset of ImageNet’s training images to serve as the shadow dataset. Additionally, we adjusted the size of each image in both the shadow and downstream datasets to dimensions of 224×224×3 as Google does in the encoder pre-training stage [5]. We set $\alpha=0.1,\beta=20$ in Algorithm 1, and $\mu=0.1$ in Algorithm 2 to scale the loss terms to the same range. Moreover, Algorithm 1 and Algorithm 2 here use a batch size of 16, and learning rates of $10^{-4}$ and 0.005 respectively. Other experiment settings are the same as those in Appendix D.6.3.
Appendix E Ethical Statements
This research is focused on identifying potential weaknesses in SSL that could be exploited by backdoor attacks. Our approach is guided by ethical standards, and we intend to enhance, not compromise, system security. We aim to promote awareness and encourage the creation of strong protective measures by demonstrating current susceptibilities within SSL frameworks.
Appendix F Limitation
The limitation of the current research is its exclusive focus on computer vision (CV) models. This specificity might not address potential vulnerabilities or behaviors applicable to other types of models, such as those used in natural language processing (NLP) or audio signal processing.
Appendix G Broader Impacts
For potential positive societal impacts, first by exposing potential vulnerabilities in SSL models, the research encourages the development of more robust security measures. This can lead to safer deployment of AI systems in critical applications such as healthcare, autonomous vehicles, and financial services, where security is paramount. Moreover, Highlighting the vulnerabilities in AI systems can spur more research into ethical AI practices. This includes developing guidelines and frameworks to ensure that AI technologies are used responsibly and do not inadvertently harm users or perpetuate biases.
For potential negative societal impacts, public knowledge of vulnerabilities in AI systems can lead to a loss of trust in these technologies. If people believe that AI systems can be easily manipulated or compromised, they may be hesitant to rely on these technologies, slowing down technological adoption and potential benefits.